OpenStudy (anonymous):

Help! Please and thank you!!:D (x^3)^3/4

OpenStudy (anonymous):

\[(x ^{3})_{4}^{3}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

express it in radical form not the answer but the way it is written above:)

OpenStudy (radar):

\[x ^{9/4}=x ^{4/4}x ^{4/4}x ^{1/4}=xxx ^{1/4}=x ^{2}x ^{1/4}=x ^{2} \sqrt[4]{x}\]

OpenStudy (anonymous):

Thank you I appreciate you taking your time to help me solve this:)!!

OpenStudy (radar):

U R welcome, I hope I have it correct.

OpenStudy (anonymous):

You were correct! Once again Thank you!!:)